### v3

# failed: mhchem

Abstract
This file contains supporting information for the paper β Topological and conventional nano-photonic waveguides for chiral integrated quantum opticsβ
Figure S1: Schematics and parameters for the waveguides used in the main text. (a) Schematic for the W1 waveguide where and . (b) Schematic for the glide-plane waveguide, where and . (c) Schematic for the zig-zag VH waveguide, where , and . (d) Location and radius of the 12 holes, closest to the interface in the initial bearded interface VH waveguide design. (e) Location and radius of the 12 holes, closest to the interface in the optimised bearded interface VH waveguide design in terms of changes relative to the original waveguide. A dielectric constant of was used throughout.
To optimise the bearded interface waveguide, an inverse design technique is used, comprised of Figures of merit (FOM) used in combination to yield more favourable properties for the topologically non-trivial mode including higher group indexes within a single mode region, and better overlap of the electric field and chiral points. A two step optimization is run to obtain the improved design. In the first step is maximized and in the second step is minimized. , is:
where is a smooth step function (equal to 0 inside a hole edge and equal to 1 more than 40Β nm from the closest hole edge), is the chiral contrast, is the Purcell factor, and is a function that encourages a certain magnitude and position of single mode bandwidth. The second phase of the optimization is focused on improving the dispersion, with a figure of merit, , of:
where is the frequency where the topologically non-trivial band first bends back, else, in the absence of a bend, it is the frequency of the topologically non-trivial band at . The parameter is the maximum frequency of the topologically non-trivial modeβs lowest frequency single mode bandwidth, is a weight to balance the third term, and is the maximum in the topologically non-trivial modeβs lowest frequency single mode bandwidth. This FOM works to eliminate any bending back of the topologically non-trivial mode and to increase the maximum single mode group index. The improved guided mode group index can be seen in the flatter band of Fig 1 (e) (ii) in comparison to (d) ii. More information about this optimisation and design of the optimised bearded interface waveguide can be found in Ref [1].
S2 Identification of Waveguide Regions
Figure S2: (a-c) Transmission of QD photoluminescence (PL) through an (a) W1, (b) glide-plane, and (c) VH waveguide, respectively. In each case, PL was generated using high-power non-resonant excitation of QDs located in a grating coupler at one end of the waveguide, and collected from the coupler at the opposite end of the device after transmission through the waveguide. The inset in (a) provides a reference PL spectrum for QDs in the bulk. The inset in (b) shows the PL signal measured from one outcoupler when exciting the QDs located in the glide-plane waveguide, revealing high finesse peaks within the slow light (and multimode) spectral window.
Considering the case of the W1 waveguide, we can see a prominent drop in transmission at 920nm, an indication of the band edge. For the glide-plane waveguide, we observed high finesse modes in the wavelength range of nm. This occurrence can be attributed to the overlapping slow light (flat band) spectral window for the two modes of the glide-plane waveguide. By identifying the regions of low and high group velocity, and locating the band edge from regions of low transmission we were able to work out the operational single mode bandwidths of the various waveguides from which we exclusively collected chiral data.
S3 Calculation of -factor
Figure S3: (a)-(e) Detailed view of dipole positions (yellow crosses) used in calculating the positional and spectral dependence of the -factor with regard to a) W1, b) glide-plane, c) zig-zag VH, d) original bearded VH and e) optimised bearded VH waveguide respectively. The irreducible repeating unit of each waveguide is shaded in blue. (f) Schematic of the FDTD -factor simulations carried out. A dipole source with in-plane circular polarisation () is surrounded by a series of transmission boxes that record the radiative power into the waveguide mode, in-plane loss, and out-of-plane loss.
The -factor of a waveguide in the contect of chiral quantum optics, quantifies the coupling efficiency of a quantum emitter into cavity/waveguide mode. It is defined as the spontaneous emission rate into guided modes over total decay rate. That is,
where are decay rates into a guided and non-guided mode respectively. The calculation of the -factor for this work is done through finite-difference-time-domain (FDTD) simulations, using the the commercial FDTD software package, Lumerical [2]. The simulations consist of measuring the power throughput of a circularly polarised dipole emitter, collected from a series of transmission/power monitors surrounding the source. The -factor is then calculated as the fraction of power coupled into the propagating modes over the total power injected from the emitter. Transmission monitors that account for out-of-plane and in-plane losses are placed two lattice constants above the waveguide and six lattice constants away from the waveguide centre respectively, while those for the waveguide modes are placed at 1Β΅m, 2Β΅m, 3Β΅m and 4Β΅m away from the dipole source for transmission comparison as well as uncertainty calculations, as is shown in Fig. S3. The simulation region is in , in and in , with layers of perfect matched layers (PML) at the x, y and z boundaries. The simulation is run with a mesh of 20 cells per lattice constant. A refractive index of and is used for the GaAs dielectric slab and air holes respectively. We vary the circularly polarised dipoleβs position within the least irreducible unit of the waveguide, up to away from the waveguide centre (see Fig.S3 for more details with regard to each waveguide).
During the -factor calculations, we observed minor, high frequency fringes in the transmission data caused by weak reflection at the PML region and the finite sampling rate inherent to the nature of the Fourier transforms of FDTD. In order to remove the fringes, we apply a SavitzkyβGolay Filter. The filter is applied to all three elements (i.e. forward, backward and total electric field power) used for the factor calculation.
S4 Loss into free space modes at chiral points
Since the -factor characterises the fraction of spontaneous emission into waveguide propagating mode, it is therefore closely associated with the Purcell-factor, which describes the spontaneous emission rate enhancement of a dipole transition by its surrounding environment in the weak coupling regime. For waveguides the Purcell factor is proportional to the group index and the electric field intensity, and can be written as: [3]
where is Bloch mode profile at position , is a unit vector that denotes dipole orientation, is group velocity, is dielectric permittivity, is the period of the waveguide. The relation between Purcell Factor and -factor is given by,
where is Purcell factor, is a factor representing the loss into free space modes, which takes a value between 0 and 1. By calculating both the -factor as described in the previous section and the Purcell enhancement of the dipole sources at chiral points, we are able to extract the loss into free space modes at chrial points for each waveguide at the frequency point indicated by the red circles in Fig.1 in the main text. If the waveguide contained multiple locations where the chiral contrast was 1, we choose the point with the highest Purcell enhancement. For each waveguide we extract values for of:
Waveguide Type
W1
Standard GP
Optimised GP
Zig-zag VH
Original bearded VH
Optimised bearded VH
S5 Spectral and spatial dependence of -factor
Figure S4: Frequency (horizontal-axis) and position (vertical-axis) dependence of -factor for a) W1, b) glide-plane, c) zig-zag VH, d) original bearded VH and e) optimised bearded VH waveguide. See Fig.S3 (a-e) for detailed dipole positions. Range of the colour-map is set to so as to highlight the contrast.
The spectral and position dependence of -factor for the five waveguides explored in the main text is shown in Fig S4. Dipole positions are labelled first along the vertical axis (upward, away from the waveguide centre), then the horizontal axis (rightward, along the propagation direction). The frequency range for each waveguide is deduced from the single-mode regime of its band structure (see Fig. 1 in the main text).
For the W1 waveguide, the single-mode region goes across the whole band, with a gradual decrease in group velocity, which reaches zero at the Brillouin Zone (BZ) edge. It has a near-invariant mode profile, contributing to a monotonic increase in the Purcell Factor against mode frequency. It is interesting to see a considerable -factor in the presence of slow-light even at positions in the waveguide that have an absence of high electric field intensity.
As for the standard glide-plane waveguide, its single-mode region only covers a small fraction of the upper band. Since there is no slow-light effect in this region and the electric field is well concentrated in the waveguide centre, the -factor is lower than the W1 waveguide. Dispersion is approximately linear in the frequency range of interest, resulting in a near-constant -factor across the single-mode region.
The Zig-Zag Topological VH waveguide is a single-mode waveguide. The modes possess a broad mode profile, with a mode waist radius of up to depending on the relative contrast of two triangular holes [4]. Although this waveguide has a single mode slow light region, this region overlaps with the bulk crystal band and so can not confine the modes within the waveguide centre (see the dark region of short wavelength modes in Fig.S4). Furthermore, the broad mode profile makes the mode less concentrated, giving rise to low -factors generally throught the waveguideβs operation.
Similar to the standard glide-plane waveguide, the original bearded VH waveguide has a slow-light region but unfortunately, once again it is multi-mode. Given that these modes tend to possess opposite polarisation, chirality is compromised and the slow-light region is no longer appropriate for chiral applications. With the additional consideration of its broad mode profile, the original bearded VH waveguide performs the worst in the -factor comparison.
A significant improvement of the optimised bearded VH waveguide is the breaking of glide-plane symmetry which resolves the multi-mode issue by shifting the high-order mode upward. The introduction of a slow-light region away from the BZ edge brings about high Purcell enhancement for topologically-protected, gap guided modes with high electric field concentration. As a result, the optimised version of the topological Bearded VH outperforms all other waveguides explored in the main text in terms of -factor .
S6 QD chirality asymmetry
Figure S5: (a) Frequency of the asymmetry in the chiral contrast measured from the left and right out-coupler for the W1 waveguide. (b) spectra of an asymmetric QD, showing the difference in contrast viewed from the right out-coupler in red and the left out-coupler in blue.
A notable characteristic observed across all waveguides, consistent with findings from other experiments and QD wafers[5, Barik2020, 6], is the presence of asymmetry in the chirality of the QDs when comparing the measured chirality from the left and right out-couplers of the waveguide. In Fig 3 (f-i), we can observe examples of QD spectra that closely align with the predictions generated by the FDTD simulations illustrated in Fig 2 (j). These spectra exhibit the anticipated behavior, where the outputs from the right out-coupler (represented in red) and the left out-coupler (represented in blue) are symmetric, acting as mirrored versions of each other. They display the same chiral contrast but with opposite intensities. However, the spectra of Fig S5 (b) presents an instance of asymmetric behavior, where the chiral contrast observed from the left out-coupler surpasses that observed from the right out-coupler. Fig S5 (a) shows the distribution of the asymmetry, showing the number of QDs that were found within the W1 waveguide with a given difference between the contrast as calculated from the left out-coupler, and the contrast as calculated from the right out-coupler. This discrepancy in behavior warrants a thorough investigation into its possible origin and the implications it holds for the practical applications of these devices.
S7 Slab Waveguide
Figure S6: (a) Top down view of 3D-FDTD simulation with 2 transmission monitors and and a 3D optical power monitor . (a) average chiral contrast and (b) average -factor with respect to the dot position (starting in the middle of the waveguide at 0nm and moving towards the edge at 140nm).
While the bulk of our work has focused on photonic crystal waveguides, for a complete comparison, we consider a simple slab waveguide. Here we analyse how the properties of chiral QD emission in a slab waveguide changes according to its position along the width of the structure. Using FDTD simulations and employing a circularly polarised dipole source, we can compare changes in average chiral contrast () and average -factor () from normalised transmission monitors placed either side of the dipole ( and ) with a 3D power monitor placed around the dipole source (layout shown in Fig S6 (a)).
With simulated results we can calculate and using:
where and is the mean transmission over the wavelength range 900-960 nm and is the ratio of the injected power into the simulation region with the power that couples into the waveguide. To ensure the best possible coupling of a single fundamental mode into the waveguide, we restrict out width nm (with a depth of 170 nm) in addition to mode expansion monitors which discard the minor coupling of higher-order modes.
From Fig S6 (c), we can see that circular dipole sources within the slab waveguide can show high contrast (max() = 92%), but the maximum -factors possible are significantly lower than that of photonic crystals.
S8 Optimised Glide-plane
Figure S7: (a) Band structure for the optimised glide-plane wavetguides showing superated upper and lower guided modes, without a multi mode region. (b) The electric field intensity at the interface of the waveguide (c) The degree of circular polarisation of the electric field at the centre of the slab () (d) -factor of the waveguide at different physics and spectral points in the waveguide.
For the main text, the simplest versions of the conventional waveguides were used, for a fair comparison to the topological waveguide designs that havenβt been optimised. However in this supplementary section we highlight a key candidate waveguide, the optimised glide-plane waveguide. Waveguides of this nature have featured predominately in work in quantum optics [7], [8], with the design that is explored here coming from reference [8].
The optimisation of this waveguide, is mostly focused on creating a favourable band structure. The and electric field map of the waveguide interface is shown in Fig S7 (b-c), which is a very similar field profile as seen in Fig 2 (b) for the un-optimsied waveguide.
In comparison to the un-optimised version, this waveguide has no multi mode region, and instead has single-mode operation from both guided modes of the waveguide. In the un-optimised waveguide, the slow-light region of operation, is also in a multi mode region. Because the two guided modes often have opposite circular polarisations of their electric fields for a single point at the interface of the waveguide, the multi mode region renders highly directional chiral emission impossible.
The slow light region of the waveguide provides a significant improvement in the -factor of the waveguide at all points within the interface. The drawbacks of the standard glide-plane waveguide design are addressed by the optimised version. The wide slow-light region at BZ edge leads to high even if the emitter is placed away from the centre. In figure S7 (d), we see a line of high -factor at Β 0.282 a/ corresponding to the slow light region.